# Modified Dai-Liao Spectral Conjugate Gradient Method with Application to Signal Processing

### v2

@@ 3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
In this section, we introduce a modified spectral conjugate gradient (CG) method, inspired by ideas from quasi-Newton methods.
@@
Following the methodology in [11], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
@@
Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then, we have the following:
@@
(a) Note that $\langle\nabla f(x_{0}),d_{0}\rangle=-\|\nabla f(x_{0})\|^{2}$. Then the inequality (21) holds for $n=0$. We show that the inequality (21) holds for $n>0$. Let $n\in\mathbb{N}_{0}$. Note that $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. From (18), (19) and (20), we get
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
@@
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
From Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}\nabla f(x_{n+1})$ and $v_{n}=\frac{\sqrt{2p}\langle\nabla f(x_{n+1}),d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
$\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\leq\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+\frac{p\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}.$ (23)
On the other hand, by Cauchy-Schwarz inequality, we have
$q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}\leq|q|\frac{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}{\|d_{n}\|^{2}}=|q|\|\nabla f(x_{n+1})\|^{2}.$ (24)
@@
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $\leq-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+|q|\|\nabla f(x_{n+1})\|^{2}$
$=-\left(\theta_{n+1}-\left(\frac{1}{4p}+|q|\right)\right)\|\nabla f(x_{n+1})\|^{2}.$
(b) Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. Then, from (21), we get $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Therefore, $\{d_{n}\}$ is a descent direction. ∎
@@
Quasi-Newton methods are a prominent class of iterative optimization techniques that do not require computing the exact Hessian matrix at each iteration. The primary motivation behind these methods is to achieve the rapid convergence rate of the classical Newton method while reducing the computational cost. Instead of calculating the Hessian matrix directly, quasi-Newton methods construct a sequence of Hessian approximations. Inspired by the quasi-Newton direction, the search direction $d_{n+1}$ defined by (20) is designed to mimic the behavior of the quasi-Newton method. Consequently, we aim to determine a suitable spectral parameter $\theta_{n+1}$ such that
$-B_{n+1}^{-1}\nabla f(x_{n+1})\approx-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n},$
where $B_{n+1}$ is an approximation of the Hessian matrix $\nabla^{2}f(x_{n+1})$. Taking inner product with $B_{n+1}s_{n}$ gives that
$-\langle s_{n},\nabla f(x_{n+1})\rangle\approx-\theta_{n+1}\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle,$
@@
$\theta_{n+1}\approx\frac{1}{\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle).$
Since $B_{n+1}$ must satisfies the modified secant condition $B_{n+1}s_{n}=z_{n}$, we have
${\theta_{n+1}}$ $=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.\right.$
@@
We neglect $\langle s_{n},\nabla f(x_{n+1})\rangle$ in (25), which leads to
@@
We denote the $\theta_{n+1}$ obtained in equation (26) by $\theta_{n+1}^{R}$ and the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{N}$. To obtain the sufficient descent condition and the bounded property of the spectral parameters, we propose the following spectral parameters:
@@
where $\eta$ and $\tau$ are positive constants. Relation (28) and (29) lead to
@@
Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that
     $10<\delta<\sigma<$ 1.
@@
Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\epsilon$, stop.
@@
$d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
      where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral
    parameter is given by (28) or (29).
@@
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-\eta\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}.$ (31)
(ii) If $\|\nabla f(x_{n})\|\neq 0$, then, from (9) and (31), we see that
@@
which, from (7), implies that $h_{n}=\nu$ and hence
@@
    From (8) and (31), we can see that $\{f(x_{n})\}$ is decreasing. Thus, the sequence $\{x_{n}\}$ generated by Algorithm 1 is contained in $C_{0}$.
@@
    From Assumption (A2), we have
@@
    $\leq 2ML+\|\nabla f(x_{0})\|:=\gamma\text{ for all }x\in C_{0}.$ (34)
@@
Since $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$. Then, there exists $\epsilon>0$ such that $\|\nabla f(x_{n})\|\geq\epsilon$ for all $n\in\mathbb{N}_{0}$. This together with (16) implies that
@@
Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3. Since $\theta_{n}>\frac{1}{4p}+|q|$, it follows from Remark 6 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Also from strong Wolfe condition (9), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$. Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 are satisfied. Thus, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}-\theta_{n+1}^{2}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}},$
where $b_{1}=\left(\frac{1}{4p}+|q|\right)\frac{1}{1+\sigma^{2}}$. It follows from (35) that for all sufficiently large $n$, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (36)
From (36) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
@@
Suppose the claim of the theorem is not true, i.e., there exists a positive constant $\xi$ such that
@@
From the Lipschitz continuity of the $\nabla f$ and (39), we obtain
$t_{n}^{p,q}$ $\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|$
$\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
Thus, the relations (19), (10), ((ii)), (38), (39), along with Cauchy-Schwarz inequality, conclude
@@
Note that $\theta_{n+1}\leq\tau$ by (30). Hence, from (20), (40), we get
@@
From (41), we observe that
@@
From (37) and Lemma 8, we have
@@
which contradicts (42). Therefore, $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0$. ∎

### v3

@@ 3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
In this section we introduce a modified spectral conjugate gradient (CG) descent method, inspired by ideas from quasi-Newton methods.
@@
Following the methodology in [KafakiGh2014], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
@@
Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then the following statements hold:
@@
(a) Note that $\langle\nabla f(x_{0}),d_{0}\rangle=-\|\nabla f(x_{0})\|^{2}$. Therefore inequality (21) holds for $n=0$. We claim that the inequality (21) holds for $n>0$. Let $n\in\mathbb{N}_{0}$. Note that $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. For simplicity, let us denote $\nabla f(x_{n+1})=F_{n+1}$. Then, using (18), (19) and (20), we get
$\langle F_{n+1},d_{n+1}\rangle$
$=\langle F_{n+1},-\theta_{n+1}F_{n+1}+\beta_{n}^{MDDL}d_{n}\rangle$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{t_{n}^{p,q}\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}.$
@@
$\langle F_{n+1},d_{n+1}\rangle$ $=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
Using Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}F_{n+1}$ and $v_{n}=\frac{\sqrt{2p}\langle F_{n+1},d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
$\frac{\langle F_{n+1},d_{n}\rangle\langle F_{n+1},z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\leq\frac{1}{4p}\|F_{n+1}\|^{2}+\frac{p\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}.$ (23)
On the other hand, by the Cauchy-Schwarz inequality, we have
$q\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\|d_{n}\|^{2}}\leq|q|\frac{\|d_{n}\|^{2}\|F_{n+1}\|^{2}}{\|d_{n}\|^{2}}=|q|\|F_{n+1}\|^{2}.$ (24)
@@
$\langle F_{n+1},d_{n+1}\rangle$ $\leq-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{1}{4p}\|F_{n+1}\|^{2}+|q|\|F_{n+1}\|^{2}$
$=-\left(\theta_{n+1}-\left(\frac{1}{4p}+|q|\right)\right)\|F_{n+1}\|^{2}.$
(b) Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. Then, employing (21), we get $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Therefore, $\{d_{n}\}$ is a descent direction as claimed. ∎
@@
Quasi-Newton methods are a prominent class of iterative optimization methods that avoid computing the exact Hessian at every step. Their main goal is to retain the fast convergence associated with Newton’s method while keeping the computational burden low. Rather than evaluating the true Hessian, these methods update an approximate one at each iteration. Motivated by the structure of quasi-Newton directions, the search direction $d_{n}$ in (20) is constructed to emulate this behavior. Our objective, therefore, is to choose an appropriate spectral parameter $\theta_{n+1}$ such that
$B_{n+1}^{-1}\nabla f(x_{n+1})\approx-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n},$
where $B_{n+1}$ is an approximation of the Hessian matrix $\nabla^{2}f(x_{n+1})$. Taking the inner product with $B_{n+1}s_{n}$, we find that
$\langle s_{n},\nabla f(x_{n+1})\rangle\approx-\theta_{n+1}\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle,$
@@
$\theta_{n+1}\approx\frac{1}{\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle).$ (25)
From (7), $B_{n+1}$ must satisfy the modified secant condition $B_{n+1}s_{n}=z_{n}$. It follows from (25) that
${\theta_{n+1}}$ $\approx\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
[…]
@@
Neglecting $\langle s_{n},\nabla f(x_{n+1})\rangle$ in (26), we arrive at
@@
We denote the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{R}$ and the $\theta_{n+1}$ obtained in equation (28) by $\theta_{n+1}^{N}$. To ensure that the sufficient-descent condition holds and to keep the spectral parameters bounded, we introduce the following choices for the spectral parameters:
@@
where $\eta$ and $\tau$ are positive constants. Relations (29) and (30) lead to
@@
Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ be such that
     $0<\delta<\sigma<1$.
@@
Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\varepsilon$, stop, where $\|u\|_{\infty}=\max_{1\leq i\leq m}|u_{i}|$ for
    $u=(u_{1},\ldots,u_{m})\in\mathbb{R}^{m}$.
@@
$d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
    where the CG parameter $\beta_{n}^{MDDL}$ is given by (19) and
      the spectral parameter $\theta_{n}$ is given by (29) or (30).
@@
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-\eta\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}_{0}.$ (32)
(ii) If $\|\nabla f(x_{n})\|\neq 0$, then using (10) and (32), we see that
@@
which, by (8), implies that $h_{n}=\nu$ and hence
@@
    Employing (9) and (32), we can see that $\{f(x_{n})\}$ is decreasing. Thus, the sequence $\{x_{n}\}$ generated by Algorithm 1 is contained in $C_{0}$.
@@
    From Assumption (A1), there exists $M>0$ such that $\|x\|\leq M$ for all $x\in C_{0}$. Then using Assumption (A2), we have
@@
    $\leq 2ML+\|\nabla f(x_{0})\|=\gamma\text{ for all }x\in C_{0}.$ (35)
@@
Since $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, there exists $\epsilon>0$ such that $\|\nabla f(x_{n})\|\geq\epsilon$ for all $n\in\mathbb{N}_{0}$. When combined with (16), this implies that
@@
Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3. Since $\theta_{n}>\frac{1}{4p}+|q|>0$, it follows from Remark 6 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Also from strong Wolfe condition (10), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$ for all $n\in\mathbb{N}_{0}$. Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 are satisfied. Thus, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}\left[B-\theta_{n+1}^{2}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}\right],$
where $B=\left(\frac{1}{4p}+|q|\right)^{2}\frac{1}{1+\sigma^{2}}$. It follows from (36) that for all sufficiently large $n$, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq B\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (37)
Using (37) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a real function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
@@
Suppose, for the sake of contradiction, that the assertion of the theorem does not hold. Then there exists a positive constant $\xi$ such that
@@
Thus, relations (19), ((ii)), (39), (40), along with the Cauchy-Schwarz inequality, yield
@@
Note that $\theta_{n+1}\leq\tau$ by (31). Hence, using (20) and (41), we arrive at
@@
Using (42), we observe that
@@
Employing (38) and Lemma 8, we find that
@@
which contradicts (43). Therefore, $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0$, as asserted. ∎